OpenStudy (anonymous):

The interior angles of a polygon equal 360°. How many sides does this polygon have?

OpenStudy (lastdaywork):

Use formula - Sum of interior angles = (n-2) * pi radian = (n-2) * 180 degree where n is the number of sides of polygon

OpenStudy (anonymous):

Thanks!

OpenStudy (lastdaywork):

:)